OpenStudy (anonymous):

I need to find the mid segment of a trapezoid the gave me one of the bases = 30 and the other base = 2+6 and they have the mid segment labeled as 3x+4, i know that it equals 1/2 the sum of the bases but i can not figure it out

OpenStudy (anonymous):

the other base is 2x+6 my bad

OpenStudy (anonymous):

3x+4=0.5(30+2+6) 3x+4=19 3x=15 x=5